OpenStudy (anonymous):

Determine the nature of the solutions of the equation x^2=X+10=0

OpenStudy (anonymous):

complex roots

OpenStudy (anonymous):

THE CHOICES ARE, ONE REAL, TWO REAL, OR TWO NON REAL SOLUTIONS

OpenStudy (anonymous):

discriminant = 1^2 - 4*1 *-10 = 1 + 40 = 41 solution is two real roots

OpenStudy (anonymous):

two non real solution

OpenStudy (anonymous):

x^2=X+10=0 it must b x^2-X+10=0

OpenStudy (anonymous):

rearrange the eqution x^2 - x - 10 = 0 a=1, b = -1 and c = -10 b^2 - 4ac = (-1)^2 - 4 x 1 x -10 which is positive number

OpenStudy (anonymous):

its confusing with the two =

OpenStudy (anonymous):

ive ignored the second one

OpenStudy (anonymous):

x^2-X+10=0 is correct

OpenStudy (anonymous):

cali - what is the correct question?

OpenStudy (anonymous):

x^2-X+10=0

OpenStudy (anonymous):

HELLO?

OpenStudy (anonymous):

two non real solution